# SafeGen: Mitigating Unsafe Content Generation in Text-to-Image Models

### v1

@@
Abstract
Text-to-image (T2I) models, such as Stable Diffusion, have exhibited remarkable performance in generating high-quality images from text descriptions in recent years. However, text-to-image models may be tricked into generating not-safe-for-work (NSFW) content, particularly in sexual scenarios. Existing countermeasures mostly focus on filtering inappropriate inputs and outputs, or suppressing improper text embeddings, which can block explicit NSFW-related content (e.g., naked or sexy) but may still be vulnerable to adversarial prompts—inputs that appear innocent but are ill-intended. In this paper, we present SafeGen, a framework to mitigate unsafe content generation by text-to-image models in a text-agnostic manner. The key idea is to eliminate unsafe visual representations from the model regardless of the text input. In this way, the text-to-image model is resistant to adversarial prompts since unsafe visual representations are obstructed from within. Extensive experiments conducted on four datasets demonstrate SafeGen’s effectiveness in mitigating unsafe content generation while preserving the high-fidelity of benign images. SafeGen outperforms eight state-of-the-art baseline methods and achieves 99.1% sexual content removal performance. Furthermore, our constructed benchmark of adversarial prompts provides a basis for future development and evaluation of anti-NSFW-generation methods.
Warnings: This paper contains sexually explicit imagery and discussions of pornography that some readers may find disturbing, distressing, and/or offensive. To mitigate the offensiveness to readers, we showcase most of the explicit images in a painting style.
1 Introduction
@@
Figure 1: Despite defending against the generation of sexually explicit images prompted by naive cues, prior methods can be bypassed or compromised by adversarial prompts. SafeGen eliminates explicit visual representations that inherently share high similarity within text-to-image (T2I) models, achieving text-agnostic mitigation against adversarial prompts since unsafe visual representations are removed from within.
@@
Various strategies have been proposed to prevent unethical image generation. Existing methods mainly prevent unsafe image generation with external [10, 11, 12] or internal [13, 14] defenses. Specifically, external defense methods employ plug-and-play safety filters to detect inappropriate textual inputs [10] or visual outputs [11] when generating images. Although external safety filters are efficient to deploy, they can be easily removed at the code level [15], rendering them ineffective in open-sourced models. Filters can also be employed to censor not-safe-for-work (NSFW) text-image paired data and retrain the Stable Diffusion 2.1 (SD-V2.1) [12] from scratch, taking as long as 200,000 hours. Internal approaches [13, 14] modify the T2I model itself. Prior internal approaches are text-dependent as they aim to instruct the T2I model to neutralize sex-related words. They require predefined NSFW concepts to steer away from the unsafe latent regions [13] or fine-tune model parameters to suppress inappropriate texts [14]. Unfortunately, as shown in Figure 1 (1st row), existing methods can be induced by adversarially-designed input to provide sexually explicit content [16, 17].
@@
To address this problem, we aim to establish a text-agnostic framework to mitigate unsafe content generation with T2I models, even in the presence of any adversarial prompts. In this paper, we analyze the difficulty encountered by prior countermeasures and propose SafeGen  [^1] SafeGen embodies dual significance: (1) Safe Generation; (2) Endowing the text-to-image model with Safe Gene., a novel text-agnostic defense method. SafeGen regulates the vision-only self-attention layers to remove the unsafe image generation capability from an already-trained T2I model, as depicted in Figure 1 (2nd row). SafeGen can also complement and seamlessly integrate with existing defense methods to further enhance the overall performance of unsafe image generation.
@@
Our first contribution is to develop an in-depth analysis of the threat of adversarial prompts. The external countermeasures, such as text- and image-based safety filters are ready-to-use and do not affect T2I models’ normal image generation. However, the under-generalization of detection models renders them inadequate in filtering adversarial prompts and unsafe images. Existing internal defenses are text-dependent, i.e., suppressing the model response to a series of sexually related words. Nevertheless, our investigation unveils their limitations in eliminating seemingly innocent but sexually connoted phrases, which are difficult to enumerate. For instance, a sample from the I2P dataset [18]—“[an adult film actor] in an orgy with [an adult film actress] and octopussy”—can make text-dependent methods ineffective. The reason is that the names of porn stars and some polysemous words carry conceptual implications of pornography. We present a detailed theoretical and experimental analysis in §3, which renders the demand for a text-agnostic paradigm.
@@
Our second contribution is to design a text-agnostic model editing technique that removes the capability of creating sexually explicit images from T2I models. Since these models are normally trained or fine-tuned on text-image paired data, designing a text-agnostic method is non-trivial. To tackle this challenge, we first trace back to the generation process of T2I models, where text-dependent and text-independent information are combined to produce the image. The text-dependent information is produced by cross-attention layers to provide textual guidance. The text-independent (i.e., vision-only) information is produced by self-attention layers to make the generated image close to the real image distribution and thus can be fine-tuned with only image samples. Therefore, we propose to modify the self-attention layers to remove sexually explicit images from the “real” image distribution utilizing a small number of image samples. In this way, we achieve lightweight and text-agnostic model modification, stopping the model from creating sexually explicit images even under sexual implications.
@@
Our third contribution is an extensive evaluation with eight baselines on a novel benchmark that comprises representative and diverse test samples. We construct prompt samples in four categories, i.e., three adversarial datasets: manually-tailored, optimization-based, and real-world picture-labeling prompts, alongside a benign COCO-25k prompt dataset. Besides the representative manually-tailored I2P dataset [18], consisting of NSFW prompts shared on lexica.art, we curate 400 optimization-based prompts containing sexually suggestive concepts by reproducing the latest attack [16]. For real-world prompts, we utilize the cutting-edge image-captioning model BLIP2 [19] to provide text that closely aligns with the semantic context of images, yielding 56,000 samples. Extensive experiments verify that SafeGen achieves the best performance in suppressing sexually explicit image generation while preserving the generation of high-fidelity benign images. We also explore the integration of SafeGen with different existing techniques, further heightening its effectiveness. We have open-sourced our implementation [20] of SafeGen to contribute to responsible AI research.
Summary of Contributions. Our theoretical and technical contributions are as follows:
  • We reveal the risk of adversarial prompts through theoretical and experimental analysis, raising an alarm about sexually explicit image generation for the community.
  • We make the first attempt to design a text-agnostic model governance technique for T2I models, suppressing sexually explicit image generation regardless of the textual input.
  • We construct a comprehensive benchmark for evaluating T2I models with adversarial and benign prompts, based on which we verify the effectiveness of our method compared with existing works.
2 Background
2.1 Diffusion Models
@@
Different from classical generative models such as Generative Adversary Network (GAN) [21] and Variational Autoencoder (VAE) [22] that synthesize images from sampled distributions in one shot, denoising diffusion models (e.g., DDPM [1], DDIM [2]) divide image generation into step-by-step sub-tasks, achieving state-of-the-art performance [23]. Apart from image generation [24], diffusion models have also been successfully applied to other modalitiy, e.g., text [25], video [26], and audio [27].
@@
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\mathtt{U}}(x_{t},t))+\sigma_{t}n,$ (1)
@@
Figure 2: Inference workflow of text-to-image Stable Diffusion. The user input is converted into embeddings and projected through cross-attention layers in each denoising step.
2.2 Text-to-Image (T2I) Generation
The success of denoising diffusion models also boosts the advancement of Text-to-Image (T2I) generative models like Stable Diffusion (SD) and Latent Diffusion [28], which gain significant attention recently. T2I models are multi-modal generation models that take texts as input, conditioned on which, visually realistic and semantically consistent images are created.
Stable Diffusion [28] is an extension to Latent Diffusion, incorporating knowledge from pre-trained CLIP [29] instead of BERT [30] as the text encoder and utilizing a more extensive training subset of LAION-5B [31]. As depicted in Figure 2, Stable Diffusion models work in a lower-dimensional latent space $z$, which speeds up the diffusion process while preserving image quality. Apart from vision-only self-attention layers in the denoising diffusion probalistic model (DDPM), Stable Diffusion models integrate additional cross-attention layers to inject embeddings of contextual input into the U-Net.
To enhance high-quality image generation that is consistent with user’s semantics and improve image diversity, T2I models [3, 5, 26] widely embrace classifier-free guidance [32, 33], which involves both a conditional and an unconditional denoising diffusion processes, i.e., $\epsilon_{\mathtt{U}}(z_{t},c,t)$ and $\epsilon_{\mathtt{U}}(z_{t},t)$, respectively. The predicted noise $\widetilde{\epsilon}_{\mathtt{U}}(z_{t},c,t)$ at time step $t$ is
$\widetilde{\epsilon}_{\mathtt{U}}(z_{t},c,t)=\epsilon_{\mathtt{U}}(z_{t},t)+\eta(\epsilon_{\mathtt{U}}(z_{t},c,t)-\epsilon_{\mathtt{U}}(z_{t},t)).$ (2)
@@ 2.2 Text-to-Image (T2I) Generation
2.3 Attention Mechanism in T2I Models
@@
Figure 3: Diagram of a cross-attention layer (in the dashed box) in text-to-image models. Text-based attention matrices $\mathbf{W_{K}}$ and $\mathbf{W_{V}}$ transform each token’s embedding into $\mathbf{K}$ and $\mathbf{V}$, respectively. Similarly, the matrix $\mathbf{W_{Q}}$ transforms visual latent to $\mathbf{Q}$.
2.3.1 Text-Dependent Cross-Attention Layers
@@
$\mathbf{M}=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{m}}\right)$ (3)
@@
Since these layers generate textual information that guides image generation, existing works [13, 14] tried to neutralize sex-related embeddings to avoid creating pornography. Nevertheless, adversarial prompts may contain implicit hints but not explicit sex-related concepts, bypassing these defenses. Discussions on existing defense methods will be detailed in §3.
2.3.2 Vision-Only Self-Attention Layers
Slightly different from cross-attention, self-attention [34] transforms the input sequence e.g., an image, into $\mathbf{Q,K,V}$ matrices and computes attention scores within itself, as depicted in Figure 4). With its superior capability of capturing intricate relationships and dependencies at pixel level, self-attention mechanism plays an important part in T2I generation [28, 5, 4], as well as other vision tasks, e.g., object detection [35], image segmentation [36], and image captioning [37].
@@ 2.3.2 Vision-Only Self-Attention Layers
Figure 4: Diagram of self-attention. The query, key, and value $\mathbf{Q,K,V}$ vectors are all obtained by the learned attention matrices $\mathbf{W_{Q},W_{K},W_{V}}$ transforming the same visual latent.
2.4 Threat Model
@@
2.4.1 Adversary
  • Objective. The adversary’s primary objective is to allure T2I models to generate unsafe content. The adversary may leverage adversarial prompts to bypass external mechanism (e.g., filter-based detection) and nullify internal techniques (e.g., explicit concepts suppression) in T2I models.
@@ 2.4.2 Model Governor
  • Objectives. The model governor has two primary objectives. The first objective is to safeguard T2I models from generating unsafe content under adversarial prompts. The second objective is to ensure high-quality image generation in response to benign prompts.
@@
  • Capabilities. The model governor has full access to the T2I model’s parameters, e.g., optimizing the whole model or editing specific module. The model governor can integrate complementary techniques, such as safe latent diffusion (SLD) [13] that aims to enhance the safety of T2I models from a textual perspective.
3 Analysis on Adversarial Prompts

### v2

# SafeGen: Mitigating Sexually Explicit Content Generation in Text-to-Image Models

@@
Abstract.
Text-to-image (T2I) models, such as Stable Diffusion, have exhibited remarkable performance in generating high-quality images from text descriptions in recent years. However, text-to-image models may be tricked into generating not-safe-for-work (NSFW) content, particularly in sexually explicit scenarios. Existing countermeasures mostly focus on filtering inappropriate inputs and outputs, or suppressing improper text embeddings, which can block sexually explicit content (e.g., naked) but may still be vulnerable to adversarial prompts—inputs that appear innocent but are ill-intended. In this paper, we present SafeGen, a framework to mitigate sexual content generation by text-to-image models in a text-agnostic manner. The key idea is to eliminate explicit visual representations from the model regardless of the text input. In this way, the text-to-image model is resistant to adversarial prompts since such unsafe visual representations are obstructed from within. Extensive experiments conducted on four datasets and large-scale user studies demonstrate SafeGen’s effectiveness in mitigating sexually explicit content generation while preserving the high-fidelity of benign images. SafeGen outperforms eight state-of-the-art baseline methods and achieves 99.4% sexual content removal performance.
Warnings: This paper contains sexually explicit imagery and discussions of pornography that some readers may find disturbing, distressing, and/or offensive.
Text-to-Image Model, Sexually Explicit, Safety, Unsafe Mitigation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3670295isbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Human and societal aspects of security and privacyccs: Theory of computation Models of computation
1. Introduction
@@
Figure 1. Despite defending against the generation of sexually explicit images prompted by naive cues, prior methods can be bypassed or compromised by adversarial prompts. SafeGen eliminates explicit visual representations that inherently share high similarity within text-to-image (T2I) models, achieving text-agnostic mitigation against adversarial prompts since unsafe visual representations are removed from within.
@@
Various strategies have been proposed to prevent unethical image generation. Existing methods mainly prevent unsafe image generation with external (text_safety_classifier, ; huggingface_safety_checker, ; SD-2-1, ) or internal (schramowski2023safe, ; gandikota2023erasing, ) defenses. Specifically, external defense methods employ plug-and-play safety filters to detect inappropriate textual inputs (text_safety_classifier, ) or visual outputs (huggingface_safety_checker, ) when generating images. Although external safety filters are efficient to deploy, they can be easily removed at the code level (reddit_remove_filter, ), rendering them ineffective in open-sourced models. Filters can also be employed to censor not-safe-for-work (NSFW) text-image paired data and retrain the Stable Diffusion 2.1 (SD-V2.1) (SD-2-1, ) from scratch, taking as long as 200,000 hours. Internal approaches (schramowski2023safe, ; gandikota2023erasing, ; gong2024reliable, ) modify the T2I model itself. Prior internal approaches are text-dependent as they aim to instruct the T2I model to neutralize sex-related words. They require predefined NSFW concepts to steer away from the unsafe latent regions (schramowski2023safe, ) or fine-tune model parameters to suppress inappropriate texts (gandikota2023erasing, ). Unfortunately, as shown in Figure 1 (1st row), existing methods can be induced by adversarially-designed input to provide sexually explicit content (yang2023sneakyprompt, ; qu2023unsafe, ).
@@
To address this problem, we aim to establish a text-agnostic framework to mitigate sexually explicit content generation with T2I models, even in the presence of any adversarial prompts. In this paper, we analyze the difficulty encountered by prior countermeasures and propose SafeGen, a novel text-agnostic defense method. SafeGen regulates the vision-only self-attention layers to remove the explicit image generation capability from an already-trained T2I model, as depicted in Figure 1 (2nd row). SafeGen can also complement and seamlessly integrate with existing defense methods to further enhance the overall performance of explicit image mitigation.
@@
Our initial step is to deeper understand the inadequacy of existing countermeasures against the threat posed by adversarial prompts. External countermeasures, such as text- and image-based safety filters are ready-to-use and do not affect T2I models’ normal image generation. However, the under-generalization of detection models renders them insufficient, as evidenced by a high detection miss rate of 23.8% in our user study, when filtering adversarial prompts and unsafe explicit images. Existing internal defenses are text-dependent, i.e., suppressing the model response to a series of sexual-related words. Nevertheless, our investigation unveils their limitations in eliminating seemingly innocent but sexually connoted phrases, which are difficult to enumerate. For instance, a sample from the I2P dataset (I2P, )—“[an adult film actor] in an orgy with [an adult film actress] and octopussy”—can make text-dependent methods ineffective. This is attributed to the names of porn stars and some polysemous words carry conceptual implications of explicitness. We review prior attacks and defenses on T2I models, and analyze the impact of adversarial prompts in §3, which renders the demand for a text-agnostic paradigm and motivates our design of SafeGen.
@@
Our first contribution is to design a text-agnostic model editing technique that removes the capability of creating sexually explicit images from T2I models. Since these models are normally trained or fine-tuned on text-image paired data, designing a text-agnostic method is non-trivial. To tackle this challenge, we first trace back to the generation process of T2I models, where text-dependent and text-independent information are combined to produce the image. The text-dependent information is produced by cross-attention layers to provide textual guidance. The text-independent (i.e., vision-only) information is produced by self-attention layers to make the generated image close to the real image distribution and thus can be fine-tuned with only image samples. Therefore, we propose to modify the self-attention layers to remove sexually explicit images from the “real” image distribution utilizing a small number of image samples. In this way, we achieve lightweight and text-agnostic model modification, stopping the model from creating sexually explicit images even under sexual implications.
@@
Our second contribution is an extensive evaluation involving multiple objective metrics and large-scale user studies, comparing eight baseline defenses on a novel benchmark that comprises representative and diverse test samples. We construct prompt samples in four categories, i.e., three adversarial datasets: manually-tailored, optimization-based, and real-world picture-labeling prompts, alongside a benign COCO-25k prompt dataset. Besides the representative manually-tailored I2P dataset (I2P, ), consisting of NSFW prompts shared on lexica.art, we curate 400 optimization-based prompts containing sexually suggestive concepts by reproducing the latest attack (yang2023sneakyprompt, ). For real-world prompts, we utilize the cutting-edge image-captioning model BLIP2 (li2023blip2, ) to provide text that closely aligns with the semantic context of images, yielding 56,000 samples. Extensive experiments verify that SafeGen achieves the best performance in suppressing sexually explicit image generation while preserving the generation of high-fidelity benign images, from both objective and human-centric perspectives. We also explore the integration of SafeGen with different existing techniques, further heightening its effectiveness. We have open-sourced our implementation (our_released_code, ) of SafeGen to contribute to responsible AI research.
Contributions. Our primary contributions are outlined below: $\bullet$ New Technique. We summarize the inadequacy of existing defenses against the generation of sexually explicit content, which motivates us to design a pioneering text-agnostic model governance technique for T2I models, termed SafeGen. Our approach identifies the importance of self-attention layers and effectively suppresses sexually explicit content generation regardless of the textual input, while maintaining high-quality benign generation with negligible false positives.
$\bullet$ New Benchmark and Findings. We construct a comprehensive benchmark for evaluating the capability of T2I models to handle both adversarial and benign prompts in terms of generating sexually explicit content. Based on this benchmark, our extensive experiments demonstrate SafeGen’s superior performance relative to eight recognized baseline defenses through objective metrics along with large-scale user studies. We also demonstrate that SafeGen can seamlessly complement existing text-based defenses, and discuss the potential of addressing over-censorship issues.
2. Background
2.1. Diffusion Models
@@
Different from classical generative models such as Generative Adversary Network (GAN) (goodfellow2014generative, ) and Variational Autoencoder (VAE) (kingma2013auto, ) that synthesize images from sampled distributions in one shot, denoising diffusion models (e.g., DDPM (ho2020denoising, ), DDIM (song2020denoising, )) divide image generation into step-by-step sub-tasks, achieving state-of-the-art (SOTA) performance (dhariwal2021diffusion, ). Apart from image generation (kawar2022denoising, ), diffusion models have also been successfully applied to other modalitiy, e.g., text (gong2022diffuseq, ), video (ho2022imagen, ), and audio (kong2020diffwave, ).
@@
(1) $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\mathtt{U}}(x_{t},t))+\sigma_{t}n,$
@@
Figure 2. Inference workflow of text-to-image Stable Diffusion. The user input is converted into embeddings and projected through cross-attention layers in each denoising step.
2.2. Text-to-Image (T2I) Generation
The success of denoising diffusion models also boosts the advancement of Text-to-Image (T2I) generative models like Stable Diffusion (SD) and Latent Diffusion (rombach2022high, ), which have gained significant attention recently. T2I models are multi-modal generation models that take texts as input, conditioned on which, visually realistic and semantically consistent images are created.
Figure 3. Diagram of a cross-attention layer (in the dashed box) in text-to-image models. Text-based attention matrices $\mathbf{W_{K}}$ and $\mathbf{W_{V}}$ transform each token’s embedding into $\mathbf{K}$ and $\mathbf{V}$, respectively. Similarly, the matrix $\mathbf{W_{Q}}$ transforms visual latent to $\mathbf{Q}$.
Stable Diffusion (rombach2022high, ) is an extension to Latent Diffusion, incorporating knowledge from pre-trained CLIP (CLIP, ) instead of BERT (devlin2018bert, ) as the text encoder and utilizing a more extensive training subset of LAION-5B (schuhmann2022laion, ). As depicted in Figure 2, Stable Diffusion models work in a lower-dimensional latent space $z$, which speeds up the diffusion process while preserving image quality. Apart from vision-only self-attention layers in the denoising diffusion probalistic model (DDPM), Stable Diffusion models integrate additional cross-attention layers to inject embeddings of contextual input into the U-Net.
To enhance high-quality image generation that is consistent with user’s semantics and improve image diversity, T2I models (SD-1-4, ; dalle2_openai, ; ho2022imagen, ) widely embrace classifier-free guidance (ho2022classifier, ; nichol2021glide, ), which involves both a conditional and an unconditional denoising diffusion processes, i.e., $\epsilon_{\mathtt{U}}(z_{t},c,t)$ and $\epsilon_{\mathtt{U}}(z_{t},t)$, respectively. The predicted noise $\widetilde{\epsilon}_{\mathtt{U}}(z_{t},c,t)$ at time step $t$ is
(2) $\widetilde{\epsilon}_{\mathtt{U}}(z_{t},c,t)=\epsilon_{\mathtt{U}}(z_{t},t)+\eta(\epsilon_{\mathtt{U}}(z_{t},c,t)-\epsilon_{\mathtt{U}}(z_{t},t)).$
@@ 2.2. Text-to-Image (T2I) Generation
2.3. Attention Mechanism in T2I Models
@@
2.3.1. Text-Dependent Cross-Attention Layers
@@
(3) $\mathbf{M}=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{m}}\right)$
@@
Since these layers generate textual information that guides image generation, existing works (schramowski2023safe, ; gandikota2023erasing, ) tried to neutralize sex-related embeddings to avoid creating pornography. Nevertheless, adversarial prompts may contain implicit hints but not explicit sex-related concepts, bypassing these defenses. Discussions on existing defense methods will be detailed in §3.
2.3.2. Vision-Only Self-Attention Layers
Slightly different from cross-attention, self-attention (Vaswani2017AttentionIA, ) transforms the input sequence e.g., an image, into $\mathbf{Q,K,V}$ matrices and computes attention scores within itself, as depicted in Figure 4). With its superior capability of capturing intricate relationships and dependencies at pixel level, self-attention mechanism plays an important part in T2I generation (rombach2022high, ; dalle2_openai, ; midjourney, ), as well as other vision tasks, e.g., object detection (Gu2020PyramidCS, ), image segmentation (Petit2021UNetTS, ), and image captioning (Guo2020NormalizedAG, ).
@@ 2.3.2. Vision-Only Self-Attention Layers
Figure 4. Diagram of self-attention. The query, key, and value $\mathbf{Q,K,V}$ vectors are all obtained by the learned attention matrices $\mathbf{W_{Q},W_{K},W_{V}}$ transforming the same visual latent.
2.4. Threat Model
@@
2.4.1. Adversary
  • $\bullet$
    Objective. The adversary’s primary objective is to allure T2I models to generate sexually explicit content. The adversary may leverage adversarial prompts to bypass external mechanism (e.g., filter-based detection) and nullify internal techniques (e.g., explicit concepts suppression) in T2I models.
  • $\bullet$
@@ 2.4.2. Model Governor
    Objectives. The model governor has two primary objectives. The first objective is to safeguard T2I models from generating explicit content under adversarial prompts. The second objective is to ensure high-quality image generation in response to benign prompts.
  • $\bullet$
@@
    Capabilities. The model governor has full access to the T2I model’s parameters, e.g., optimizing the whole model or editing specific module. The model governor can integrate complementary techniques, such as safe latent diffusion (SLD) (schramowski2023safe, ) that aims to enhance the safety of T2I models from a textual perspective.
3. Related Work & Motivation